# GS-SDF: LiDAR-Augmented Gaussian Splatting and Neural SDF for Geometrically Consistent Rendering and Reconstruction

### v1

@@ IV-C2 FAST-LIVO2 Datasets
We compare the qualitativer rendering results in a free trajectory (Drive), as shown in Fig. 8, where our method outperforms the others in retaining detailed textures of the floor with the help of geometric regularization. It is noted that both the LiDAR-augmented NeRF-based method (M2Mapping) and 3DGS-based method (Ours) give physically grounded rendering results, while the NeRF-based method shows more complete rendering at the view lacking observations and 3DGS-based method shows more detailed rendering overall. We also compared the reconstruction results over the diverse types of methods, as shown in Fig. 7. It is shown that LiDAR-based methods (VDBFusion, SHINE-Mapping, M2Mapping, Ours) offer far more detailed geometry than the vision-only methods (2DGS, PGSR).
@@ IV-D1 Initialization
In this section, we only apply the render regularization (Eq. 5) as a pure image supervision to address the importance of the proposed Gaussian initializations (Sec. III-C) in converging a good structure of Gaussians. As shown in Fig. 9, the original initialilzation, with random sampling position from the input point clouds and random attributes initialization, falls into unreasonable structures (Fig. 9(b)). While, the proposed SDF-aided physically-grounded initialization gives a structure that more in line with the real scenarios. A color initialization assigns a proper appearance to the Gaussians, which is essential to stabilize the structure of Gaussians. As shown in Fig. 10, the inited Gaussian splats with proper structure but wrong color are deviated during early training.
@@ IV-D2 Geometric regularization
We compare the geometric regularization methods between the render regularization [9], center regularization and shape regularization (Sec. III-D2), as shown in Fig. 8(g)-8(i). The render regularization (PSNR: 27.24, SSIM: 0.805, Fig. 8(g)) improves the geometrically consistency to a extent, but still produces false structure on the flat surfaces under the scenes with less multi-view constrains. The center regularization (PSNR: 28.60, SSIM: 0.861, Fig. 8(h)) shows its limited improvement in aligning Gaussians with the SDF field, still producing blur rendering on the ground. And our proposed shape regularization (PSNR: 28.66, SSIM: 0.866, Fig. 8(i)) shows better alignment with the surface to retain the details of the bricks. As shown in Fig. 9(e)’s rendered normal image, the shape regularization gives a more detailed and physically grounded structure of Gaussians for shaper rendering results.
@@ IV-E Efficiency Analysis
The training time for the Replica Room-2 scene takes 16.6 minutes and the average rendering time for a 1200x680 image takes about 9.9 milliseconds and reaches 101.1 frames per second (FPS). We evaluate the efficiency of the compared NeRF-based method, M2Mapping, which takes 53.9 milliseconds for rendering a 1200x680 image (18.4 FPS). As our method trains both the NSDF and 3DGS, it takes more training time than the other 3DGS-based methods, like 2DGS (9.1 minutes / 103.2 FPS) but the inference time is comparable to the other methods and shows a significant improvement over the NeRF-based methods.

### v2

@@ IV-C2 FAST-LIVO2 Datasets
We compare the qualitative rendering results in a free trajectory (Drive), as shown in Fig. 8, where our method outperforms the others in retaining detailed textures of the floor with the help of geometric regularization. It is noted that both the LiDAR-augmented NeRF-based method (M2Mapping) and 3DGS-based method (Ours) give physically grounded rendering results, while the NeRF-based method shows more complete rendering at views lacking observations, and the 3DGS-based methods show more detailed rendering overall. We also compared the reconstruction results over the diverse types of methods, as shown in Fig. 7. It is shown that LiDAR-based methods (VDBFusion, SHINE-Mapping, M2Mapping, Ours) offer far more detailed geometry than the vision-only methods (2DGS, PGSR).
@@ IV-D1 Initialization
In this section, we only apply the render regularization (Eq. 5) as a pure image supervision to address the importance of the proposed Gaussian initializations (Sec. III-C) in converging to a good structure of Gaussians. As shown in Fig. 9, the original initialization, with random sampling position from the input point clouds and random attributes initialization, falls into unreasonable structures (Fig. 9(b)). The proposed SDF-aided physically-grounded initialization gives a structure that is more in line with the real scenarios. A color initialization assigns a proper appearance to the Gaussians, which is essential to stabilize the structure of the Gaussians. As shown in Fig. 10, the initialized Gaussian splats with proper structure but wrong color are deviated during early training.
@@ IV-D2 Geometric regularization
We compare the geometric regularization methods between the render regularization [9], center regularization and shape regularization (Sec. III-D2), as shown in Fig. 8(g)-8(i). The render regularization (PSNR: 27.24, SSIM: 0.805, Fig. 8(g)) improves the geometric consistency to an extent, but still produces false structure on the flat surfaces under scenes with less multi-view constraints. The center regularization (PSNR: 28.60, SSIM: 0.861, Fig. 8(h)) shows its limited improvement in aligning Gaussians with the SDF field, still producing blur rendering on the ground. And our proposed shape regularization (PSNR: 28.66, SSIM: 0.866, Fig. 8(i)) shows better alignment with the surface to retain the details of the bricks. As shown in Fig. 9(e)’s rendered normal image, the shape regularization gives a more detailed and physically grounded structure of Gaussians for sharper rendering results, while the evaluation metrics do not show a significant improvement.
@@ IV-E Efficiency Analysis
The training time for the Replica Room-2 scene takes 16.6 minutes, and the average rendering time for a 1200x680 image takes about 9.9 milliseconds and reaches 101.1 frames per second (FPS). We evaluate the efficiency of the compared NeRF-based method, M2Mapping, which takes 53.9 milliseconds for rendering a 1200x680 image (18.4 FPS). As our method trains both the NSDF and 3DGS, it takes more training time than the other 3DGS-based methods, like 2DGS (9.1 minutes / 103.2 FPS), but the inference time is comparable to the other methods and shows a significant improvement over the NeRF-based methods.